# Sampling from the Mean-Field Stationary Distribution

### v2

@@ 3 Technical Ingredients
Our general approach for sampling from the stationary distribution $\pi$ in either (1.1) or (2.2) is to directly apply an off-the-shelf sampler for the finite-particle stationary distribution $\mu^{1:N}$. The theoretical guarantees for this procedure require two main ingredients: (1) control of the “bias”—i.e., the error incurred by approximating $\pi$ by the $1$-particle marginal of $\mu^{1:N}$—and (2) verification of isoperimetric properties which allow for fast sampling from the measure $\mu^{1:N}$.
@@ 3.1.1 Pairwise McKean–Vlasov Setting
Remark. Note that from (1.1), we typically would expect $C_{\mathsf{LSI}}^{2}(\pi)$ to also scale as $\sigma^{4}$ (e.g., in the case when $V$ and $W$ are $\alpha$-uniformly convex for $\alpha>0$). Therefore, Assumption 3 is typically invariant to the scaling of $\sigma$ and can be satisfied even for $\sigma\searrow 0$. Under these assumptions, we obtain the following result via a modification of the argument of [LL23]. Compared to their work, we remove the assumption of boundedness of $\lVert\nabla W\rVert$, and our argument is simpler since we apply it directly to the stationary distribution.
@@
The following weaker result consists of two parts. The first, a Wasserstein propagation of chaos result, is based on [Szn91]. The second, building on the first, is a uniform-in-time entropic propagation of chaos bound following from a Fisher information bound. The arguments are similar to those in [Mal01, Mal03], albeit simplified (since we work at stationarity) and presented here with explicit constants.
@@ 3.1.2 General McKean–Vlasov Setting
In the more general case where we aim to minimize ($\mathsf{gE}$) for a generic functional $\mathcal{F}$ of the form $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$, we impose the following assumptions. They can be largely seen as generalizations of the conditions for the pairwise case, and they are inherited from [CRW22, SWN23]. There is an additional convexity condition (Assumption 5), which in the pairwise McKean–Vlasov setting amounts to positive semidefiniteness of the kernel $(x,y)\mapsto W(x-y)$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$; thus, in general, the following assumptions are incomparable with the ones in §3.1.1.
@@
The functional $\mathcal{F}$ is smooth in the sense that for all $x,y\in\mathbb{R}^{d}$, $\nu,\nu^{\prime}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, there is a uniform constant $\beta$ such that
$\lVert\nabla_{\mathcal{W}_{2}}\mathcal{F}(\nu,x)-\nabla_{\mathcal{W}_{2}}\mathcal{F}(\nu^{\prime},y)\rVert\leq\beta\,(\lVert x-y\rVert+\mathcal{W}_{1}(\nu,\nu^{\prime}))\,.$
@@
The proximal Gibbs measures satisfy (LSI) with a uniform constant: namely, it holds that $C_{\mathsf{LSI}}(\pi)\vee\sup_{\mu\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})}C_{\mathsf{LSI}}(\pi_{\mu})\leq\overline{C}_{\mathsf{LSI}}$.
Remark. These assumptions taken together cover settings not covered in the preceding sections, including optimization of two-layer neural networks. See [CRW22, Remark 3.1] and §4.2.
Under these assumptions, we can derive an entropic propagation of chaos bound by following the proof of [CRW22]. Through a tighter analysis, we are able to reduce the dependence on the condition number $\kappa\coloneqq\overline{C}_{\mathsf{LSI}}\beta/\sigma^{2}$ from $\kappa^{2}$ to $\kappa$.
@@
Among these assumptions, the hardest to verify is the uniform LSI of Assumption 7. Following [SWN23], we introduce the following sufficient condition for the validity of Assumption 7; see Lemma 21 for a more precise statement.
@@
Assumptions 6 and 8 imply Assumption 7 with an explicit constant $\overline{C}_{\mathsf{LSI}}$, given in terms of $d$, $B$, $\beta$, $\lambda$, and $\sigma$.
@@ 3.2 Isoperimetric Properties of the Stationary Distributions
In this section, we verify the isoperimetric properties of $\pi,\mu^{1:N}$, with proofs provided in §B.
@@ 3.2.1 Pairwise McKean–Vlasov Setting
If $V$, $W$ satisfy Assumptions 1 and 4 (i.e., $V$ and $W$ have bounded Hessians), then the potential for (1.1), i.e., $\log(1/\pi)$, is $\frac{2}{\sigma^{2}}\,(\alpha_{V}+\alpha_{W})$-convex and $\frac{2}{\sigma^{2}}\,(\beta_{V}+\beta_{W})$-smooth. By the Bakry–Émery condition, $\pi$ satisfies (LSI) with parameter $C_{\mathsf{LSI}}(\pi)\leq\nicefrac{{\sigma^{2}}}{{2\,(\alpha_{V}+\alpha_{W})}}$.
@@
If $V$ and $W$ satisfy Assumption 4, then $\log(1/\mu^{1:N})$ is $\frac{2}{\sigma^{2}}\,(\alpha_{V}+\frac{N}{N-1}\,{\alpha_{W}^{-}})$-convex. [^2] Only the negative part of $\alpha_{W}$ contributes to the strong log-concavity of $\mu^{1:N}$. This is consistent with [Vil03, Theorem 5.15], which asserts that when $\alpha_{W}>0$, the interaction energy $\mu\mapsto\iint W(x-y)\,\mu(\mathrm{d}x)\,\mu(\mathrm{d}y)$ is $\alpha_{W}$-strongly displacement convex over the subspace of probability measures with fixed mean, but only weakly convex over the full Wasserstein space.
@@
In this section, we provide an easily verifiable condition, based on the Holley–Stroock condition [HS87], for this assumption to hold with an $N$-independent constant.
@@
The potentials $V$ and $W$ can be decomposed as $V=V_{0}+V_{1}$ and $W=W_{0}+W_{1}$ such $V_{0}$, $W_{0}$ satisfy Assumption 4 and $\operatorname{\mathsf{osc}}(V_{1}),\operatorname{\mathsf{osc}}(W_{1})<\infty$, where for a function $U:\mathbb{R}^{d}\to\mathbb{R}$ we define $\operatorname{\mathsf{osc}}(U)\coloneqq\sup U-\inf U$.
Under this assumption, a careful application of the Holley–Stroock perturbation principle yields the following lemma.
@@
$C_{\mathsf{LSI}}(\pi)$ $\leq\frac{\sigma^{2}}{2\,(\alpha_{V_{0}}+\alpha_{W_{0}})}\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\operatorname{\mathsf{osc}}(V_{1})+\operatorname{\mathsf{osc}}(W_{1})\bigr{)}\Bigr{)}\,$
$C_{\mathsf{LSI}}(\mu^{1:N})$ $\leq\frac{\sigma^{2}}{2\,(\alpha_{V_{0}}+\frac{N}{N-1}\,{\alpha_{W_{0}}^{-}})}\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\operatorname{\mathsf{osc}}(V_{1})+\operatorname{\mathsf{osc}}(W_{1})\bigr{)}\Bigr{)}\,.$
@@ 3.2.2 General McKean–Vlasov Setting
In the setting ($\mathsf{gE}$) with $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$, we verify that Assumption 8 yields (LSI) for $\mu^{1:N}$ with an $N$-independent constant. See Lemma 22 for a more precise statement.
@@
In the mean-field Langevin setting of §2.1.2, suppose that Assumption 8 holds. Then, Assumption 9 holds with $C_{\mathsf{LSI}}(\mu^{1:N})$ depending on $d$, $B$, $\lambda$, and $\sigma$, but not on $N$.
@@ 4 Sampling from the Mean-Field Target
In this section, we present results for sampling from $\pi$. As outlined in Algorithm 1, we use off-the-shelf log-concave samplers to sample from $\mu^{1:N}$, during which we access the first-order [^3] For our results involving the proximal sampler, we also assume access to a proximal oracle for simplicity. oracle for $\mu^{1:N}$ (i.e., an oracle for evaluation of $\log\mu^{1:N}$ up to an additive constant, and for evaluation of $\nabla\log\mu^{1:N}$). For $N$ sufficiently large, the first particle given by Algorithm 1 is approximately distributed according to $\pi$: for $\hat{\mu}^{1:N}$ the law of the output of the log-concave sampler and its $1$-particle marginal distribution $\hat{\mu}^{1}$,
@@
${}^{+}$ $\nicefrac{{\sqrt{\alpha}}}{{\sigma}}\,\mathcal{W}_{2}$ 134 ${\kappa d^{1/3}}/{\varepsilon^{2/3}}$ ${d^{1/2}}/{\varepsilon}$
@@
The algorithms in the table refer to: Langevin Monte Carlo (LMC); underdamped Langevin Monte Carlo (ULMC); discretizations of the underdamped Langevin diffusion via the randomized midpoint method [SL19] or the shifted ODE method [FLO21] (${}^{+}$); and implementation of the proximal sampler [LST21, Che+22] via the Metropolis-adjusted Langevin algorithm or via ULMC (MALA–PS and ULMC–PS respectively). Note that LMC applied to sample from $\mu^{1:N}$ is simply the Euler–Maruyama discretization of ($\mathsf{pMV}_{N}$), and likewise ULMC is the algorithm considered in [FW23]. We refer to §E for proofs and references.
To streamline the rates, we simplify the notation by defining $\beta=\beta_{V}+\beta_{W}$ if Assumption 1 holds, otherwise we use the value from Assumption 6. We let $\alpha=\alpha_{V}+{\alpha_{W}^{-}}$ under Assumption 4, $\alpha=\nicefrac{{\sigma^{2}}}{{2\max\{C_{\mathsf{LSI}}(\mu^{1:N}),C_{\mathsf{LSI}}(\pi)\}}}$ under Assumptions 2 and 9, and $\alpha=\nicefrac{{\sigma^{2}}}{{2\max\{C_{\mathsf{LSI}}(\mu^{1:N}),\overline{C}_{\mathsf{LSI}}\}}}$ in the general McKean–Vlasov setting. Finally, we let $\kappa\coloneqq\beta/\alpha$ denote the condition number. The additional assumption $\kappa\leq\sqrt{d}/\varepsilon$ will be used to simplify some of the rates.
@@ 4.1 Pairwise McKean–Vlasov Setting
Consider the strongly convex case where $\alpha=\alpha_{V}+{\alpha_{W}^{-}}>0$. The prior work [BS23] also considered the problem of sampling from the mean-field stationary distribution $\pi$, with $\sigma^{2}=2$. If we count the number of calls to a gradient oracle for $V$, their complexity bound reads $\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{5/3}d^{4/3}}}{{\varepsilon^{8/3}}})$ to achieve $\nicefrac{{\sqrt{\alpha}}}{{\sigma}}\,\mathcal{W}_{1}(\hat{\mu}^{1},\pi)\leq\varepsilon$. We note that their assumptions are not strictly comparable to ours. They require the interaction $W$ to be sufficiently weak, in the sense that $\beta_{W}\lesssim\alpha$, which is similar [^4] See eq. (2.24) therein; note that they have a scaling factor of $\varepsilon$ in front of their interaction term, so that our parameter $\beta_{W}$ is equivalent to their $\varepsilon\tilde{L}$. to our Assumption 3; on the other hand, they only assume $\alpha_{V}>0$, rather than $\alpha_{V}+{\alpha_{W}^{-}}>0$. Nevertheless, we attempt to make some comparisons with their work below.
Without Assumption 3, ${}^{+}$ achieves $\nicefrac{{\sqrt{\alpha}}}{{\sigma}}\,\mathcal{W}_{2}(\hat{\mu}^{1},\pi)\leq\varepsilon$ with complexity $\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{3}d^{4/3}}}{{\varepsilon^{8/3}}})$, which matches the guarantee of [BS23] up to the dependence on $\kappa$. We can also obtain guarantees in $\sqrt{\operatorname{\mathsf{KL}}}$, at the cost of an extra factor of $\kappa^{2}$.
With Assumption 3, MALA–PS has complexity $\widetilde{\mathcal{O}}(\nicefrac{{\kappa d^{5/4}}}{{\varepsilon^{3/2}}})$ and ${}^{+}$ has complexity $\widetilde{\mathcal{O}}(\nicefrac{{\kappa d^{5/6}}}{{\varepsilon^{5/3}}})$, which improve substantially upon [BS23].
To summarize, in the strongly convex case, we have obtained numerous improvements: (i) we can obtain results even without the weak interaction condition (Assumption 3); (ii) when we assume the weak interaction condition, we obtain improved complexities; (iii) our results hold in stronger metrics; (iv) our approach is generic, allowing for the consideration of numerous different samplers without needing to establish new propagation of chaos results (by way of comparison, [BS23] developed a tailored propagation of chaos argument for their non-linear Hamiltonian Monte Carlo algorithm).
@@
Both the results of [BS23] as well as our own allow for non-convex potentials, albeit under different assumptions—[BS23] require strong convexity at infinity, whereas we require (LSI) for the stationary measures $\mu^{1:N}$ and $\pi$. In order to obtain sampling guarantees with low complexity, it is important for the LSI constant of $\mu^{1:N}$ to be independent of $N$. We have provided a sufficient condition for this to hold: $V$ and $W$ are bounded perturbations of $V_{0}$ and $W_{0}$ respectively, where $\alpha_{V_{0}}+{\alpha_{W_{0}}^{-}}>0$; see Lemma 8.

### v4

@@ 3 Technical Ingredients
Our general approach for sampling from the stationary distribution $\pi$ in either (1.1) or (2.2) is to directly apply an off-the-shelf sampler for the finite-particle stationary distribution $\mu^{1:N}$. The theoretical guarantees for this procedure require two main ingredients: (1) control of the “bias”—i.e. the error incurred by approximating $\pi$ by the $1$-particle marginal of $\mu^{1:N}$—and (2) verification of isoperimetric properties which allow for fast sampling from the measure $\mu^{1:N}$.
@@ 3.1.1 Pairwise McKean–Vlasov Setting
Remark. Note that from (1.1), we typically would expect $C_{\mathsf{LSI}}^{2}(\pi)$ to also scale as $\sigma^{4}$ (e.g., in the case when $V$ and $W$ are $\alpha$-uniformly convex for $\alpha>0$). Therefore, Assumption 3 is typically invariant to the scaling of $\sigma$ and can be satisfied even for $\sigma\searrow 0$. Under these assumptions, we obtain a sharp propagation of chaos result via a similar argument as [lacker2022quantitative, lacker2023sharp]. We note that the former is more permissive regarding the constant in Assumption 3 as compared to this work.
@@
The following weaker result consists of two parts. The first, a Wasserstein propagation of chaos result, is based on [sznitman1991topics]. The second, building on the first, is a uniform-in-time entropic propagation of chaos bound following from a Fisher information bound. The arguments are similar to those in [malrieu2001logarithmic, malrieu2003convergence], albeit simplified (since we work at stationarity) and presented here with explicit constants.
@@ 3.1.2 General McKean–Vlasov Setting
In the more general case where we aim to minimize ($\mathsf{gE}$) for a generic functional $\mathcal{F}$ of the form $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$, we impose the following assumptions. They can be largely seen as generalizations of the conditions for the pairwise case, and they are inherited from [chen2022uniform, suzuki2023convergence]. There is an additional convexity condition (Assumption 5), which in the pairwise McKean–Vlasov setting amounts to positive semidefiniteness of the kernel $(x,y)\mapsto W(x-y)$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$; thus, in general, the following assumptions are incomparable with the ones in §3.1.1.
@@
The functional $\mathcal{F}_{0}$ is smooth in the sense that for all $x,y\in\mathbb{R}^{d}$, $\nu,\nu^{\prime}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, there is a uniform constant $\beta$ such that
$\lVert\nabla_{\mathcal{W}_{2}}\mathcal{F}_{0}(\nu,x)-\nabla_{\mathcal{W}_{2}}\mathcal{F}_{0}(\nu^{\prime},y)\rVert\leq\beta\,(\lVert x-y\rVert+\mathcal{W}_{1}(\nu,\nu^{\prime}))\,.$
@@
The proximal Gibbs measures satisfy (LSI) with a uniform constant: namely, it holds that $C_{\mathsf{LSI}}(\pi)\vee\sup_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}C_{\mathsf{LSI}}(\pi_{\mu})\leq\overline{C}_{\mathsf{LSI}}$.
Remark. These assumptions taken together cover settings not covered in the preceding sections, including optimization of two-layer neural networks. See [chen2022uniform, Remark 3.1] and §4.2.
Under these assumptions, we can derive an entropic propagation of chaos bound by following the proof of [chen2022uniform]. Through a tighter analysis, we are able to reduce the dependence on the condition number $\kappa\coloneqq\overline{C}_{\mathsf{LSI}}\beta/\sigma^{2}$ from $\kappa^{2}$ to $\kappa$.
@@
Among these assumptions, the hardest to verify is the uniform LSI of Assumption 7. Following [suzuki2023convergence], we introduce the following sufficient condition for the validity of Assumption 7; see Lemma 22 for a more precise statement.
@@
Assumptions 6 and 8 imply Assumption 7 with an explicit constant $\overline{C}_{\mathsf{LSI}}$, given in terms of $B$, $\beta$, $\lambda$, and $\sigma$.
@@ 3.2 Isoperimetric Properties of the Stationary Distributions
In this section, we verify the isoperimetric properties of $\pi,\mu^{1:N}$ in the ($\mathsf{pMV}$) setting, with proofs provided in §B.
@@ 3.2.1 Pairwise McKean–Vlasov Setting
If $V$, $W$ satisfy Assumptions 1 and 4 (i.e. $V$ and $W$ have bounded Hessians), then the potential for (1.1), i.e. $\log(1/\pi)$, is $\frac{2}{\sigma^{2}}\,(\alpha_{V}+\alpha_{W})$-convex and $\frac{2}{\sigma^{2}}\,(\beta_{V}+\beta_{W})$-smooth. By the Bakry–Émery condition, $\pi$ satisfies (LSI) with parameter $C_{\mathsf{LSI}}(\pi)\leq\nicefrac{{\sigma^{2}}}{{2\,(\alpha_{V}+\alpha_{W})}}$.
@@
If $V$ and $W$ satisfy Assumption 4, then $\log(1/\mu^{1:N})$ is $\frac{2}{\sigma^{2}}\,(\alpha_{V}+\frac{N}{N-1}\,{\alpha_{W}^{-}})$-convex. [^2] Only the negative part of $\alpha_{W}$ contributes to the strong log-concavity of $\mu^{1:N}$. This is consistent with [villani2003topics, Theorem 5.15], which asserts that when $\alpha_{W}>0$, the interaction energy $\mu\mapsto\iint W(x-y)\,\mu(\mathrm{d}x)\,\mu(\mathrm{d}y)$ is $\alpha_{W}$-strongly displacement convex over the subspace of probability measures with fixed mean, but only weakly convex over the full Wasserstein space.
@@
In this section, we provide an easily verifiable condition, combining a Holley–Stroock condition [HolStr1987LSI] with a weak interaction condition, for this assumption to hold with an $N$-independent constant.
@@
The potentials $V$ and $W$ can be decomposed as $V=V_{0}+V_{1}$ and $W=W_{0}+W_{1}$ such $V_{0}$, $W_{0}$ satisfy Assumption 4 and $\operatorname{\mathsf{osc}}(V_{1}),\operatorname{\mathsf{osc}}(W_{1})<\infty$, where for a function $U:\mathbb{R}^{d}\to\mathbb{R}$ we define $\operatorname{\mathsf{osc}}(U)\coloneqq\sup U-\inf U$. Furthermore, the following weak interaction condition holds:
$\frac{\sigma^{2}}{\beta_{W}\overline{C}_{\mathsf{LSI}}}\geq\sqrt{6}\,,\qquad\text{where}\qquad\overline{C}_{\mathsf{LSI}}\coloneqq\frac{\sigma^{2}}{\alpha_{V_{0}}+\frac{N}{N-1}\,{\alpha_{W_{0}}^{-}}}\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\operatorname{\mathsf{osc}}(V_{1})+\operatorname{\mathsf{osc}}(W_{1})\bigr{)}\Bigr{)}\,.$
A careful application of the Holley–Stroock perturbation principle yields the following lemma.
@@
$C_{\mathsf{LSI}}(\pi)$ $\leq\frac{\sigma^{2}}{2\,(\alpha_{V_{0}}+\alpha_{W_{0}})}\exp\Bigl{(}\frac{2}{\sigma^{2}}\,\bigl{(}\operatorname{\mathsf{osc}}(V_{1})+\operatorname{\mathsf{osc}}(W_{1})\bigr{)}\Bigr{)}\leq\frac{1}{2}\,\overline{C}_{\mathsf{LSI}}\,,$ (3.4)
$C_{\mathsf{LSI}}(\mu^{1:N})$ $\leq\overline{C}_{\mathsf{LSI}}\,.$
In particular, Assumption 3 holds.
@@ 3.2.2 General McKean–Vlasov Setting
In the setting ($\mathsf{gE}$) with $\mathcal{F}(\mu)=\mathcal{F}_{0}(\mu)+\frac{\lambda}{2}\int\lVert\cdot\rVert^{2}\,\mathrm{d}\mu$, we verify that Assumption 8 yields (LSI) for $\pi$. See Corollary 25 for a more precise statement.
@@
In the mean-field Langevin setting of §2.1.2, suppose that Assumption 8 holds. Then, Assumption 9 holds with $C_{\mathsf{LSI}}(\mu^{1:N})$ depending on $d$, $B$, $\beta$, $\lambda$, and $\sigma$, but not on $N$.
Obtaining Lemma 9 is not straightforward, and we rely on a novel argument combining heat flow estimates from [BriPed24HeatFlow] with a generalized version of the propagation of chaos result in Theorem 5; see §B.3 for details.
@@ 4 Sampling from the Mean-Field Target
In this section, we present results for sampling from $\pi$. As outlined in Algorithm 1, we use off-the-shelf log-concave samplers to sample from $\mu^{1:N}$, during which we access the first-order [^3] For our results involving the proximal sampler, we also assume access to a proximal oracle for simplicity. oracle for $\mu^{1:N}$ (i.e. an oracle for evaluation of $\log\mu^{1:N}$ up to an additive constant, and for evaluation of $\nabla\log\mu^{1:N}$). For $N$ sufficiently large, the first particle given by Algorithm 1 is approximately distributed according to $\pi$: for $\hat{\mu}^{1:N}$ the law of the output of the log-concave sampler and its $1$-particle marginal distribution $\hat{\mu}^{1}$,
@@
ULMC+ $\nicefrac{{\sqrt{\alpha}}}{{\sigma}}\,\mathcal{W}_{2}$ 134 ${\kappa d^{1/3}}/{\varepsilon^{2/3}}$ ${d^{1/2}}/{\varepsilon}$
@@
The algorithms in the table refer to: Langevin Monte Carlo (LMC); underdamped Langevin Monte Carlo (ULMC); discretizations of the underdamped Langevin diffusion via the randomized midpoint method [shen2019randomized] or the shifted ODE method [FosLyoObe21ShiftedODE] (ULMC+); and implementation of the proximal sampler [LeeSheTia21RGO, chen22improved] via the Metropolis-adjusted Langevin algorithm or via ULMC (MALA–PS and ULMC–PS respectively). Note that LMC applied to sample from $\mu^{1:N}$ is simply the Euler–Maruyama discretization of ($\mathsf{pMV}_{N}$), and likewise ULMC is the algorithm considered in [fu2023mean]. We refer to §E for proofs and references.
To streamline the rates, we simplify the notation by defining $\beta=\beta_{V}+\beta_{W}$ if Assumption 1 holds, otherwise we use the value from Assumption 6. We let $\alpha=\alpha_{V}+{\alpha_{W}^{-}}$ under Assumption 4, $\alpha=\nicefrac{{\sigma^{2}}}{{2\max\{C_{\mathsf{LSI}}(\mu^{1:N}),C_{\mathsf{LSI}}(\pi)\}}}$ under Assumptions 2 and 9, and $\alpha=\nicefrac{{\sigma^{2}}}{{2\max\{C_{\mathsf{LSI}}(\mu^{1:N}),\overline{C}_{\mathsf{LSI}}\}}}$ in the general McKean–Vlasov setting.
Finally, we let $\kappa\coloneqq\beta/\alpha$ denote the condition number. We briefly justify this terminology. If the target and all proximal Gibbs measures are strongly convex with parameter $\alpha/\sigma^{2}$, then the Bakry–Émery condition implies that $\overline{C}_{\mathsf{LSI}}\leq\tfrac{\sigma^{2}}{\alpha}$. Hence, the scale-invariant ratio $\overline{C}_{\mathsf{LSI}}\beta/\sigma^{2}$ reduces to the classical condition number $\beta/\alpha$, the ratio of the largest to smallest eigenvalues of the Hessian matrices for $V$ and $W$. Therefore, $\overline{C}_{\mathsf{LSI}}\beta/\sigma^{2}$ is a generalization of the condition number to settings beyond uniform strong convexity which allows us to state more interpretable bounds. The additional assumption $\kappa\leq\sqrt{d}/\varepsilon$ will be used to simplify some of the rates.
@@ 4.1 Pairwise McKean–Vlasov Setting
Consider the strongly convex case where $\alpha=\alpha_{V}+{\alpha_{W}^{-}}>0$. The prior work [bou2023nonlinear] also considered the problem of sampling from the mean-field stationary distribution $\pi$, with $\sigma^{2}=2$. If we count the number of calls to a gradient oracle for $V$, their complexity bound reads $\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{5/3}d^{4/3}}}{{\varepsilon^{8/3}}})$ to achieve $\nicefrac{{\sqrt{\alpha}}}{{\sigma}}\,\mathcal{W}_{1}(\hat{\mu}^{1},\pi)\leq\varepsilon$. We note that their assumptions are not strictly comparable to ours. They require the interaction $W$ to be sufficiently weak, in the sense that $\beta_{W}\lesssim\alpha$, which is similar [^4] See eq. (2.24) therein; note that they have a scaling factor of $\varepsilon$ in front of their interaction term, so that our parameter $\beta_{W}$ is equivalent to their $\varepsilon\tilde{L}$. to our Assumption 3; on the other hand, they only assume $\alpha_{V}>0$, rather than $\alpha_{V}+{\alpha_{W}^{-}}>0$. Nevertheless, we attempt to make some comparisons with their work below.
Without Assumption 3, ULMC+ achieves $\nicefrac{{\sqrt{\alpha}}}{{\sigma}}\,\mathcal{W}_{2}(\hat{\mu}^{1},\pi)\leq\varepsilon$ with complexity $\widetilde{\mathcal{O}}(\nicefrac{{\kappa^{3}d^{4/3}}}{{\varepsilon^{8/3}}})$, which matches the guarantee of [bou2023nonlinear] up to the dependence on $\kappa$. We can also obtain guarantees in $\sqrt{\operatorname{\mathsf{KL}}}$, at the cost of an extra factor of $\kappa^{2}$.
With Assumption 3, MALA–PS has complexity $\widetilde{\mathcal{O}}(\nicefrac{{\kappa d^{5/4}}}{{\varepsilon^{3/2}}})$ and ULMC+ has complexity $\widetilde{\mathcal{O}}(\nicefrac{{\kappa d^{5/6}}}{{\varepsilon^{5/3}}})$, which improve substantially upon [bou2023nonlinear].
To summarize, in the strongly convex case, we have obtained numerous improvements: (i) we can obtain results even without the weak interaction condition (Assumption 3); (ii) when we assume the weak interaction condition, we obtain improved complexities; (iii) our results hold in stronger metrics; (iv) our approach is generic, allowing for the consideration of numerous different samplers without needing to establish new propagation of chaos results (by way of comparison, [bou2023nonlinear] developed a tailored propagation of chaos argument for their non-linear Hamiltonian Monte Carlo algorithm).
@@
Both the results of [bou2023nonlinear] as well as our own allow for non-convex potentials, albeit under different assumptions—[bou2023nonlinear] require strong convexity at infinity, whereas we require (LSI) for the stationary measures $\mu^{1:N}$ and $\pi$. In order to obtain sampling guarantees with low complexity, it is important for the LSI constant of $\mu^{1:N}$ to be independent of $N$. We have provided a sufficient condition for this to hold: $V$ and $W$ are bounded perturbations of $V_{0}$ and $W_{0}$ respectively, where $\alpha_{V_{0}}+{\alpha_{W_{0}}^{-}}>0$; see Lemma 8.